# CAOTE: KV Cache Eviction for LLMs via Attention Output Error-Based Token Selection

### v3

@@ 3.1 CAOTE Score
Using Eq. (19) CAOTE scores (or eviction error) for each token can be computed in parallel as the dependency on only on attention scores and value vectors. Note that this is the firsts formulation seamlessly integrating attention scores and value vectors into a single score. Any norm can be used for computing CAOTE score and based on empirical results we choose -norm.
@@ 3.2 CAOTE with general score-based eviction methods
We consider H2O (Zhang et al., 2024b), where the scores () are based on the sum of previous attention scores, leading to during generation-phase as proved in Theorem 24 in Section B.1. In this case, simply dividing each token score by the sum of all scores maps the scores to the range and ensures that new scores follow .
@@ 3.3 Fast CAOTE Computation
We also propose a compute-efficient version of CAOTE, FastCAOTE, with negligible performance degradation and reduced computation by an order of , where is the hidden dimension of the model. Here, the pre-eviction attention output () is replaced with mean of values while everything else remains same, that is, CAOTE score for token is:
@@ 4.2 LongBench
We present the accuracy of Llama 3.1-8B-Instruct, Llama 3.2-3B-Instruct and, Qwen 2.5-3B-Instruct, Qwen 2.5-7B-Instruct using baseline eviction methods with budget of k, k, both with and without CAOTE in Table 2 and Table 3. We observe that the best performance is given by SnapKV-FastCAOTE for the Llama3 models, while for Qwen 2.5 models SnapKV-CAOTE performs the best. H2O shows improvement with CAOTE, while TOVA, SnapKV also show overall improvements, making their average accuracy closer to dense accuracy. Additional results for the k, k budget are shown in Table 7, Table 8 for Llama3 and Qwen 2.5 respectively, in Section C.1, which follow a trend similar to the k, k budgets.
@@ 4.3 Perplexity
We use the Booksum dataset (Kryściński et al., 2021) to measure generation perplexity of different eviction methods for various budgets. In Table 4, we show perplexity gap between a model using a given eviction strategy and that of the model without token eviction with cache budgets of k, k and k. We observe that when CAOTE is applied to existing eviction methods, the perplexity either improves or surpasses the perplexity of the baseline model. TOVA-FastCAOTE, SnapKV-CAOTE, and SnapKV-FastCAOTE perform best for k, k, k budgets, respectively, for Llama 3.1-8B-Instruct; for Llama 3.2-3B-Instruct, TOVA-FastCAOTE performs best with k and k budgets and SnapKV-FastCAOTE beats other methods using k and k. Perplexity results for Qwen 2.5 models are shown in Table 9 in Section C.2.
@@ 4.4 Needle In A HayStack
Lastly, we run extensive experiments on Needle-In-A-Haystack benchmark (Liu et al., 2024; Kamradt, 2023) and show quantitative results in Table 5 and visualizations for k budget Llama 3.1-8B model in Figure 2. We observe in Table 5 that H2O-FastCAOTE performs best for all budgets with Llama 3.2-3B. When using a budget k with Llama 3.1-8B, CAOTE boosted H2O outperforms TOVA, SnapKV as well. H2O-CAOTE performs best for Llama 3.1-8B with budget = {k, k} and H2O-FastCAOTE performs best fork budget for Llama3.1-8b. The gains in precision are especially high for the k budget for the Llama 3.1-8B model, with an increase of up to . We can see in Figure 2 that CAOTE improves the state-of-the-art eviction method and is able to predict beyond their budget constraints. Results for Qwen 2.5 models are shown in Table 10 in Section C.3.
4.5 Ablations
Difference in Attention Outputs
In Figure 3, we check the average change in attention score in different layers for various eviction methods to provide insight into the performance gain realized by CAOTE. We measure the difference between the dense attention-output and eviction method based attention output, both with and without CAOTE. We repeat this for each layer and average the result to verify that CAOTE results in smaller distance to dense attention-output. Figure 3 shows the normalized mean squared error between the dense attention output and eviction-based attention output for H2O, TOVA and SnapKV, both with and without CAOTE in dashed-lines and solid-lines respectively. We observe that the CAOTE variant always produces a smaller gap from dense attention output (no eviction) compared to baselines, supporting our claim that CAOTE evicts the token which minimizes deviation from true attention output. Furthermore, we show that discrepancy in attention output due to eviction (eviction error) causes discrepancy in logits in B.2; to emphasize that optimizing for eviction error results in less discrepancy in logits and, therefore, less deviation on downstream task accuracy.
Different Prompt Block Sizes
| block-size | 32 | 128 | 32 | 128 | 32 | 128 | 32 | 128 | 32 | 128 | 32 | 128 | ||||||
@@ Appendix A Limitation
Appendix B Additional Proofs
B.1 H2O scores
Theorem B.1.
@@ B.1 H2O scores
B.2 Relation of CAOTE score (eviction error) to logits
@@
Appendix C Additional Results
C.1 LongBench

### v4

@@ 3.1 CAOTE Score
Using Eq. (19) CAOTE scores (or eviction error) for each token can be computed in parallel as the dependency on only on attention scores and value vectors. Note that this is the firsts formulation seamlessly integrating attention scores and value vectors into a single score. Any norm can be used for computing CAOTE score and based on empirical results we choose -norm. Evicting multiple tokens using CAOTE formulation is discussed in Appendix B.
@@ 3.2 CAOTE with general score-based eviction methods
We consider H2O (Zhang et al., 2024b), where the scores () are based on the sum of previous attention scores, leading to during generation-phase as proved in Theorem 36 in Section D.1. In this case, simply dividing each token score by the sum of all scores maps the scores to the range and ensures that new scores follow .
@@ 3.3 Fast CAOTE Computation
We also propose a compute-efficient version of CAOTE, FastCAOTE, with negligible performance degradation and minimal compute overhead relative to CAOTE. Here, the pre-eviction attention output () is replaced with mean of values while everything else remains same, that is, CAOTE score for token is:
We find empirically, that FastCAOTE is highly correlated to CAOTE by observing high () Spearman’s correlation between them for each layer as shown in Table 8.
We provide theoretical inference time latency for using CAOTE and FastCAOTE in Appendix C, where we observe that FastCAOTE leads to minimal increase in latency.
@@ 4.2 LongBench
We present the accuracy of Llama 3.1-8B-Instruct, Llama 3.2-3B-Instruct and, Qwen 2.5-3B-Instruct, Qwen 2.5-7B-Instruct using baseline eviction methods with budget of k, k, both with and without CAOTE in Table 2 and Table 3. We observe that the best average performance is given by SnapKV-FastCAOTE for the Llama3 models, while for Qwen 2.5 models SnapKV-CAOTE performs the best. H2O shows improvement with CAOTE, while TOVA, SnapKV also show overall improvements, making their average accuracy closer to dense accuracy. Additionally, for some QA tasks (Qasper, MF-en, Musique, 2WikiMQA), H2O-(Fast)CAOTE performs best. Additional results for the k, k budget are shown in Table 9, Table 10 for Llama3 and Qwen 2.5 respectively, in Section E.1, which follow a trend similar to the k, k budgets.
@@ 4.3 Perplexity
We use the Booksum dataset (Kryściński et al., 2021) to measure generation perplexity of different eviction methods for various budgets. In Table 4, we show perplexity gap between a model using a given eviction strategy and that of the model without token eviction with cache budgets of k, k and k. We observe that when CAOTE is applied to existing eviction methods, the perplexity either improves or surpasses the perplexity of the baseline model. TOVA-FastCAOTE, SnapKV-CAOTE, and SnapKV-FastCAOTE perform best for k, k, k budgets, respectively, for Llama 3.1-8B-Instruct; for Llama 3.2-3B-Instruct, TOVA-FastCAOTE performs best with k and k budgets and SnapKV-FastCAOTE beats other methods using k and k. Perplexity results for Qwen 2.5 models are shown in Table 11 in Section E.2.
@@ 4.4 Needle In A HayStack
Lastly, we run extensive experiments on Needle-In-A-Haystack benchmark (Liu et al., 2024; Kamradt, 2023) and show quantitative results in Table 5 and visualizations for k budget Llama 3.1-8B model in Figure 2. We observe in Table 5 that H2O-FastCAOTE performs best for all budgets with Llama 3.2-3B. When using a budget k with Llama 3.1-8B, CAOTE boosted H2O outperforms TOVA, SnapKV as well. H2O-CAOTE performs best for Llama 3.1-8B with budget = {k, k} and H2O-FastCAOTE performs best fork budget for Llama3.1-8b. The gains in precision are especially high for the k budget for the Llama 3.1-8B model, with an increase of up to . We can see in Figure 2 that CAOTE improves the state-of-the-art eviction method and is able to predict beyond their budget constraints. Results for Qwen 2.5 models are shown in Table 12 in Section E.3.
@@ Appendix A Limitation
Appendix B Extending CAOTE to Multi-token Eviction
We start with simple example, without loss of generality consider evicting two tokens in the order . From (8) single token eviction error for token is
After evicting token , the updated attention output becomes:
Now, evicting token from updated attention output above yields
This leads to the following insights:
The updated attention output can be expressed as
(28) -
Eviction error of removing token after removing token is
(29) -
Substituting the expressions above, we obtain a closed-form joint eviction score:
This formulation generalizes to evicting tokens jointly (assuming first tokens here without loss of generality)
This expression highlights the combinatorial nature of multi-token eviction: for tokens with -token eviction, there are possible combinations. This makes exact computation intractable for large , motivating the need for approximate or greedy strategies.
Importantly, existing token eviction methods typically rely only on attention scores and assume independence between tokens. As a result, the relative ranking of tokens remains unchanged even after evictions, which limits their effectiveness in multi-token settings.
In contrast, CAOTE explicitly models the interplay between attention scores and value vectors, capturing how evicting one token affects the contribution of others. This introduces interdependencies between tokens during eviction, which are critical for accurate multi-token decisions.
Appendix C Inference Latency
We analyze the computational overhead of CAOTE and FastCAOTE during both prefill and generation phases. The overhead is minimal, especially for FastCAOTE.
Let be sequence length, the hidden dimension, the KV hidden dimension, the intermediate dimension, the vocabulary size, and the number of hidden layers. The total floating-point Operation count are as follows:
We compute the relative overhead for Llama3.1-8B model with different sequence lengths and show the ratio between prefill (and generation) flops to CAOTE flops in Table 6, 7
| Sequence Length | Ratio | |
| with CAOTE | with FastCAOTE | |
| 4k | 1.6e-4 | 8.9e-5 |
| 8k | 9.25e-5 | 5.28e-5 |
| 32k | 6.45e-5 | 3.69e-5 |
| Sequence Length | Ratio | |
| with CAOTE | with FastCAOTE | |
| 512 | 0.08 | 0.046 |
| 1024 | 0.15 | 0.087 |
Appendix D Additional Proofs
D.1 H2O scores
Theorem D.1.
@@ D.1 H2O scores
D.2 Relation of CAOTE score (eviction error) to logits
@@
Appendix E Additional Results
| Layer | 1 | 2 | 3 | 4 | 5 | 6 | 7 | 8 | 9 | 10 | 11 | 12 | 13 | 14 |
| 0.99 | 0.98 | 0.91 | 0.97 | 0.92 | 0.88 | 0.85 | 0.82 | 0.81 | 0.90 | 0.84 | 0.92 | 0.83 | 0.86 |
| layer | 15 | 16 | 17 | 18 | 19 | 20 | 21 | 22 | 23 | 24 | 25 | 26 | 27 | 28 |
| 0.91 | 0.88 | 0.94 | 0.96 | 0.92 | 0.95 | 0.99 | 0.99 | 0.96 | 0.98 | 0.98 | 0.88 | 0.99 | 0.96 |
E.1 LongBench
@@ E.3 Needle in a Haystack
E.4 Ablations
Difference in Attention Outputs
In Figure 3, we check the average change in attention score in different layers for various eviction methods to provide insight into the performance gain realized by CAOTE. We measure the difference between the dense attention-output and eviction method based attention output, both with and without CAOTE. We repeat this for each layer and average the result to verify that CAOTE results in smaller distance to dense attention-output. Figure 3 shows the normalized mean squared error between the dense attention output and eviction-based attention output for H2O, TOVA and SnapKV, both with and without CAOTE in dashed-lines and solid-lines respectively. We observe that the CAOTE variant always produces a smaller gap from dense attention output (no eviction) compared to baselines, supporting our claim that CAOTE evicts the token which minimizes deviation from true attention output. Furthermore, we show that discrepancy in attention output due to eviction (eviction error) causes discrepancy in logits in D.2; to emphasize that optimizing for eviction error results in less discrepancy in logits and, therefore, less deviation on downstream task accuracy.
Different Prompt Block Sizes
| block-size | 32 | 128 | 32 | 128 | 32 | 128 | 32 | 128 | 32 | 128 | 32 | 128 | ||||||